the perimeter of a rectangle is 66m if the width were double and the length were increased by 22m, the perimeter would be 136 what is the length and width of the rectangle
Answer by mangopeeler07(462)

the perimeter of a rectangle is 66m if the width were double and the length were increased by 22m, the perimeter would be 136 what is the length and width of the rectangle

2w+2L=66
2(2w)+2(L+22)=136

Solve for 2w in the first equation
2w=66-2L

Plug that into the second equation
2(66-2L)+2(L+22)=136

Distribute
132-4L+2L+44=136

Combine like terms
176-2L=136

Subtract 176 from both sides
-2L=-40

Divide both sides by -2
L=20

2w=66-2L
2w=66-2(20)
2w=66-40
2w=26
w=13

Answer:
Length=20m
Width=13m